Abc is a triangle, with D as midpoint on bc. It is given that angle adb=45 and angle acd=30. Find angle bad

Grade:9

2 Answers

Aarush Rathore
5 years ago
IN TRIANGLE ADC
 
ADB=45(EXTERIOR ANGLE)
SO ADB=DAC+ACD
DAC=ADB-ACD=45-30=15
SINCE BC IS MIDPOINT OF BC
BD=DC AND ANGLE OPPOSITE TO THEM ARE EQUAL
SO  BAD=DAC=15
Bhavuk
5 years ago
The above anwer is wrong......'.........
You can visit here for correct answer......
The above answer is wrong..................